OpenStudy (co1byjack):

How do I simplify fractions with exponents, such as 5^6/5^2?

OpenStudy (calculusxy):

remember this \(\huge \frac{a^y}{a^x} = a^{y-x}\)

OpenStudy (kendricklamar2014):

\[\frac{ 5^6 }{ 5^2} = \frac{ 5\times5\times5\times5\times5\times5 }{ 5\times5 }\]

OpenStudy (co1byjack):

Thanks Guys, KendrickLamar I thought so.

OpenStudy (kendricklamar2014):

\[= 5\times5\times5\times5\] \[= 5^4\]

OpenStudy (calculusxy):

since this question also has the same base we can do \(\huge \frac{5^6}{5^2} = 5^{6-2} = 5^x\)

OpenStudy (kendricklamar2014):

Yeah and since 5^4 is the same as 5*5*5*5, the answer would be 625. Do you agree @Co1byJack

OpenStudy (calculusxy):

and \(x\) as you may know, is 4 \(\text{Final Answer} = \huge 5^4\)

OpenStudy (co1byjack):

@KendrickLamar2014 Yes

OpenStudy (kendricklamar2014):

:)